For the fission reaction 232U + n -----> 137Te + 97Zr + 2n

(a) Calculate the amount of energy produced per mol; (b) The heat of combustion of TNT, C7H5N3O6, is 3406 kJ/mol. FInd the mass of TNT needed to produce the same energy as 1.000 mol of the fission reaction above. (c) Calculate the energy released in (a) per gram of 235 U.
Answer #1

(a) The mass converted into energy is difference between sum of the mass of reactants and sum of mass of products. Here, its

Mass of Reactant: 235.04 + 1 = 236.04 g/mol

Mass of Product: 136.9 + 96.9 + 2 = 235.8 g/mol

Difference = 236.04 - 235.8 = 0.24 g/mol

As per Einstein Equation, E = mc2

where, m is the difference in mass of the reactants consumed and products formed, in kg

c is the velocity of light in m/s = 3 X 108 m/s

E is the energy released in Joule/mol

putting the values in the equation we get

E = 0.24 X 10-3 X (3 X 108)2 X 6.023 x 1023 = 13 X 1036 J/mol or 13 X 1033 KJ/mol (Answer)

factor used for conversion of g to kg 10-3

factor used for no of molecules in a mol 6.023 x 1023